# Radiative decays of the $\Omega(2012)$ as a hadronic molecule

### v1

@@ Abstract
We present a theoretical investigation of the radiative decay process $\Omega(2012)\to\gamma\Omega$, where the $\Omega(2012)$ resonance with spin-parity $J^{P}=\frac{3}{2}^{-}$, is treated as a dynamically generated state from $\bar{K}\Xi(1530)$ and $\eta\Omega$ in $s$-wave and $\bar{K}\Xi$ in $d$-wave. The radiative decay width of the $\Omega(2012)$ is calculated using a triangular loop mechanism, where the $\Omega(2012)$ couples to the $\bar{K}\Xi(1530)$ channel. Subsequently, the final state interactions between $\Xi(1530)$ and $\bar{K}$ transition to a photon and $\Omega$ through the exchange of a $\Xi$ baryon. Our calculations yield a radiative decay width of $13.2^{+4.5}_{-3.9}$ KeV, with uncertainties arising from the model parameters. This result provides valuable insights into the nature of the $\Omega(2012)$ resonance and its decay dynamics. It is expected that the calculations presented here could be verified by future experiments, which would open a new door for studying the still elusive nature of the $\Omega(2012)$.
@@ I Introduction
On the theoretical side, an $\Omega$ excited state with a mass around 2 GeV has been predicted by many theoretical works [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18] before the $\Omega(2012)$ resonance was discovered. In particular, Ref. [8] reported a $J^{P}=3/2^{-}$ $\Omega^{*}$ resonance with a mass of about 2020 MeV within a quark model incorporating ingredients motivated by quantum chromodynamics. This state was found to decay into the $\bar{K}\Xi$ channel. Moreover, the radiative decay width of that $J^{P}=3/2^{-}$ $\Omega^{*}$ resonance with a mass of about 2020 MeV was calculated in Ref. [19] within the nonrelativistic potential model, yielding a value of $18$ keV. It is worth noting that the $\Omega^{*}$ sates with $J^{P}=1/2^{-}$ and $J^{P}=3/2^{-}$ in Refs. [8, 19] are degenerate, sharing the same mass and partial decay width to the $\gamma\Omega$ channel.
@@ II Formalism and ingredients
Figure 3: The radiative decay of $\Omega(2012)\to\gamma\Omega$ through triangle loop mechanism with the strong coupling of the $\Omega(2012)$ to the $\bar{K}\Xi(1530)$ channel.
@@ II.1 The molecular nature of $\Omega(2012)$ and its coupling to $\bar{K}\Xi(1530)$
$V_{\bar{K}\Xi^{*}\to\bar{K}\Xi^{*}}$ $=$ $V_{\eta\Omega\to\eta\Omega}=V_{\bar{K}\Xi\to\bar{K}\Xi}=0,$ (1)
$V_{\bar{K}\Xi^{*}\to\eta\Omega}$ $=$ $V_{\eta\Omega\to\bar{K}\Xi^{*}}=-\frac{3}{4f^{2}_{\pi}}(E_{K}+E_{\eta}),$ (2)
$V_{\bar{K}\Xi^{*}\to\bar{K}\Xi}$ $=$ $V_{\bar{K}\Xi\to\bar{K}\Xi^{*}}=\alpha q^{2}_{\Xi},$ (3)
@@
where $G$ is the loop function for each channel, and it can be regularized with a cutoff prescription. We refer to Ref. [38] for additional details regarding these model parameters.
@@
From Table 1, one can see that the obtained resonance parameters, mass and width, of the $\Omega(2012)$ states are in agreement with the experimental results within uncertainties. Furthermore, the larger value of the ratio ${\cal R}^{\bar{K}\pi\Xi}_{\bar{K}\Xi}$ can be also accommodated. These results again support the molecular picture of the $\Omega(2012)$ state.
Table 1: Model parameters and pole position $(M_{R},\Gamma_{R})$ of the $\Omega(2012)$ and its couplings to the involving channels.
$q_{\rm max}$ $\alpha$ $\beta$ $M_{R},\Gamma_{R}]$ $\tilde{g}_{\bar{K}\Xi^{*}}$ $\tilde{g}_{\eta\Omega}$ $\tilde{g}_{\bar{K}\Xi}$ ${\cal R}^{\bar{K}\pi\Xi}_{\bar{K}\Xi}$
(MeV) ($\times 10^{-8}\,{\rm MeV}^{-3}$) ($\times 10^{-8}\,{\rm MeV}^{-3}$) (MeV)
800 175-175 316 [2014.1, 3.4] 2.1 1.9 0.21 0.76
850 174-174 371 [2014.1, 3.4] 2.1 1.6 0.20 0.81
900 157-157 372 [2014.1, 3.4] 2.1 1.4 0.20 0.81
1000 120-120 328 [2014.2, 3.4] 2.1 1.1 0.21 0.81
1100 91-91 273 [2014.2, 3.4] 2.1 0.88 0.20 0.81
1200 69-69 219 [2014.2, 3.5] 2.1 0.73 0.21 0.78
The coupling of the $\Omega(2012)$ to each relevant channel $\tilde{g}_{i}$ ($i=\bar{K}\Xi^{*}$, $\eta\Omega$, $\bar{K}\Xi$) can be determined via the residues of the scattering amplitude $T_{ij}$ at the pole position, which reads,
$T_{ij}=\frac{\tilde{g}_{i}\tilde{g}_{j}}{\sqrt{s}-z_{R}},$ (9)
where $z_{R}=M_{R}-i\Gamma_{R}/2$ denotes the $\Omega(2012)$ pole position. The theoretical results for these coupling constants are also shown in Table 1 for each scenario of model parameters and $q_{\rm max}$. Considering the dominant radiative decay mechanism shown in Fig. 3, we need the coupling constant of $\tilde{g}_{\bar{K}\Xi^{*}}$, and in the numerical study we take
$g_{\Omega^{*}K^{-}\Xi(1530)^{0}}=g_{\Omega^{*}\bar{K}^{0}\Xi(1530)^{-}}=-\frac{1}{\sqrt{2}}\tilde{g}_{\bar{K}\Xi^{*}}.$ (10)
@@ II.2 The effective Lagrangians for the $\Xi(1530)\Xi\gamma$ and $\Omega\bar{K}\Xi$ interaction
$\mathcal{L}_{\Xi^{*}\Xi\gamma}=$ $i\frac{eg_{1}}{2m_{N}}F^{\mu\nu}\bar{\Xi}\gamma_{\mu}\gamma_{5}\Xi^{*}_{\nu}$ (11)
$-\frac{eg_{2}}{(2m_{N})^{2}}F^{\mu\nu}(\partial_{\mu}\bar{\Xi})\gamma_{5}\Xi^{*}_{\nu}+H.c.\,$,
@@
$\mathcal{L}_{\Omega\bar{K}\Xi}=\frac{f_{\Omega\Xi K}}{m_{K}}\bar{\Omega}_{v}\Xi\partial^{v}\bar{K}\,$, (12)
@@ II.3 The radiative decay $\Omega(2012)\to\gamma\Omega$
$\mathcal{M}_{\Omega(2012)\to\gamma\Omega}$ $=\int\frac{d^{4}q}{(2\pi)^{4}}\frac{\widetilde{A}F^{2}(q_{ex})}{q^{2}-m_{\Xi^{*}}^{2}+i\varepsilon}\times$ (13)
@@
where $P$, $q$, and $q_{ex}$ are the four-momenta of the $\Omega(2012)$, $\Xi(1530)$, and the exchanged hyperon $\Xi$, respectively, and
@@
$\left.+\frac{1}{3m_{\Xi^{*}}}(\gamma_{\beta}q_{\nu}-\gamma_{\nu}q_{\beta})\right],$
$A_{3}^{\nu}$ $=$ $\frac{\tilde{g}_{\bar{K}\Xi^{*}}}{\sqrt{2}}u^{\nu}(P,s_{\Omega^{*}}).$ (19)
Here, $k_{1}$ ($\lambda$) and $k_{2}$ ($s_{\Omega}$) are the four-momenta (spin polarization) of the final photon and $\Omega$ baryon. $u^{\mu}(P,s_{\Omega^{*}})$ is a Rarita-Schwinger spinor for the $\Omega(2012)$ resonance with momentum $P$ and spin polarization $s_{\Omega^{*}}$. The expression for the photon polarization vector $\varepsilon^{\mu}(k_{1},\lambda)$ are given by
@@
$\begin{gathered}P=(M_{\Omega^{*}},0,0,0),\\ q=(q_{0},\left\lvert\vec{q}\right\rvert\sin\theta\cos\phi,\left\lvert\vec{q}\right\rvert\sin\theta\sin\phi,\left\lvert\vec{q}\right\rvert\cos\theta),\\ k_{1}=(E_{\gamma},0,0,-E_{\gamma}),k_{2}=(E_{\Omega},0,0,E_{\gamma}),\end{gathered}$ (27)
@@
$E_{\gamma/\Omega}=\frac{M_{\Omega^{*}}^{2}\mp m_{\Omega}^{2}}{2M_{\Omega^{*}}},$ (28)
where $M_{\Omega^{*}}$ is the mass of $\Omega(2012)$, $E_{\gamma}(E_{\Omega})$ is the energy of the final photon ($\Omega$). We first integrate out the $q_{0}$ component utilizing the residue theorem [65] and introduce a cutoff $\Lambda$ for the integration over $|\vec{q}|$,
@@
$D$ $=$ $8\omega_{\Xi^{*}}\omega_{K}\omega_{\Xi}(P_{0}+\omega_{\Xi^{*}}+\omega_{K})(P_{0}-\omega_{\Xi^{*}}+\omega_{K}+i\epsilon)$ (30)
$\times(P_{0}-\omega_{\Xi^{*}}-\omega_{K}+i\epsilon)(E_{\gamma}+\omega_{\Xi^{*}}+\omega_{\Xi})$
$\times(E_{\gamma}-\omega_{\Xi^{*}}+\omega_{\Xi}+i\epsilon)(E_{\gamma}-\omega_{\Xi^{*}}-\omega_{\Xi}+i\epsilon)$
@@
$N_{1}$ $=$ $4\omega_{K}\omega_{\Xi}(P_{0}+\omega_{\Xi^{*}}+\omega_{K})(E_{\gamma}+\omega_{\Xi^{*}}+\omega_{\Xi})$
@@
$\times(E_{\Omega}-\omega_{K}-\omega_{\Xi})\widetilde{t}|_{\omega_{\Xi^{*}}},$
$N_{2}$ $=$ $4\omega_{\Xi^{*}}\omega_{\Xi}(P_{0}-\omega_{\Xi^{*}}+\omega_{K})(E_{\gamma}+\omega_{\Xi^{*}}+\omega_{\Xi})$
$\times(E_{\gamma}-\omega_{\Xi^{*}}+\omega_{\Xi})(E_{\gamma}-\omega_{\Xi^{*}}-\omega_{\Xi})$
@@
$N_{3}$ $=$ $4\omega_{\Xi^{*}}\omega_{K}(P_{0}+\omega_{\Xi^{*}}+\omega_{K})(P_{0}-\omega_{\Xi^{*}}+\omega_{K})$
$\times(P_{0}-\omega_{\Xi^{*}}-\omega_{K})(E_{\gamma}-\omega_{\Xi^{*}}-\omega_{\Xi})$
@@
where $P_{0}=M_{\Omega^{*}}$, $\widetilde{t}|_{x}$ is the value of $\widetilde{t}$ obtained at $q_{0}=x$ with $\widetilde{t}=\widetilde{A}F^{2}(q_{ex})$, and
$\omega_{\Xi^{*}}$ $=\sqrt{|\vec{q}|^{2}+m_{\Xi^{*}}^{2}},~~~~~\omega_{K}=\sqrt{|\vec{q}|^{2}+m_{K}^{2}},$ (32)
@@
$\Gamma_{\Omega(2012)\to\gamma\Omega}=\frac{E_{\gamma}}{32\pi M_{\Omega^{*}}^{2}}\overline{\sum_{s_{\Omega^{*}}}}\sum_{\lambda,s_{\Omega}}|\mathcal{M}_{\Omega(2012)\to\gamma\Omega}|^{2},$ (33)

### v2

@@ Abstract
We present a theoretical investigation of the radiative decay process $\Omega(2012)\to\gamma\Omega$, where the $\Omega(2012)$ resonance with spin-parity $J^{P}=\frac{3}{2}^{-}$, is treated as a dynamically generated state from $\bar{K}\Xi(1530)$ and $\eta\Omega$ in $s$-wave and $\bar{K}\Xi$ in $d$-wave. The radiative decay width of the $\Omega(2012)$ is calculated using a triangular loop mechanism, where the $\Omega(2012)$ couples to the $\bar{K}\Xi(1530)$ channel. Subsequently, the final state interactions between $\Xi(1530)$ and $\bar{K}$ transition to a photon and $\Omega$ through the exchange of a $\Xi$ baryon. Our calculations yield a radiative decay width of $13.2^{+4.5}_{-5.6}$ keV, with uncertainties arising from the model parameters. This result provides valuable insights into the nature of the $\Omega(2012)$ resonance and its decay dynamics. It is expected that the calculations presented here could be verified by future experiments, which would open a new door for studying the still elusive nature of the $\Omega(2012)$.
@@ I Introduction
On the theoretical side, an $\Omega$ excited state with a mass around 2 GeV has been predicted by many theoretical works [8, 7, 40, 30, 42, 4, 10, 9, 24, 45, 62] before the $\Omega(2012)$ resonance was discovered. In particular, Ref. [8] reported a $J^{P}=3/2^{-}$ excited $\Omega$ resonance with a mass of about 2020 MeV within a quark model incorporating ingredients motivated by quantum chromodynamics. This state was found to decay into the $\bar{K}\Xi$ channel. Moreover, the radiative decay width of that $J^{P}=3/2^{-}$ excited $\Omega$ resonance with a mass of about 2020 MeV was calculated in Ref. [23] within the nonrelativistic potential model, yielding a value of $18$ keV. It is worth noting that the excited $\Omega$ sates with $J^{P}=1/2^{-}$ and $J^{P}=3/2^{-}$ in Refs. [8, 23] are degenerate, sharing the same mass and partial decay width to the $\gamma\Omega$ channel.
@@ II Formalism and ingredients
[…]
@@ II.1 The molecular nature of $\Omega(2012)$ and its coupling to $\bar{K}\Xi(1530)$
$V_{\bar{K}\Xi(1530)\to\bar{K}\Xi(1530)}=V_{\eta\Omega\to\eta\Omega}=V_{\bar{K}\Xi\to\bar{K}\Xi}=0,$ (1)
$V_{\bar{K}\Xi(1530)\to\eta\Omega}=V_{\eta\Omega\to\bar{K}\Xi(1530)}=-\frac{3}{4f^{2}_{\pi}}(E_{K}+E_{\eta}),$ (2)
$V_{\bar{K}\Xi(1530)\to\bar{K}\Xi}=V_{\bar{K}\Xi\to\bar{K}\Xi(1530)}=\alpha q^{2}_{\Xi},$ (3)
@@
where $G$ is the loop function for each channel, and it can be regularized with a cutoff prescription. Typically, the cutoff parameter ($q_{\rm max}$) in the loop functions for the $\bar{K}\Xi(1530)$, $\eta\Omega$, and $\bar{K}\Xi$ channels are different. In this work, we adopt a common value $q_{\rm max}$ for all channels, which will be determined from the Belle collaboration experimental data. We refer to Ref. [31] for additional details regarding these model parameters.
@@
From Table 1, one can see that the obtained resonance parameters, mass and width, of the $\Omega(2012)$ states are in agreement with the experimental results within uncertainties.
Table 1: Model parameters and pole position $(M_{R},\Gamma_{R})$ (in MeV) of the $\Omega(2012)$ and its couplings to the involving channels. Parameter $q_{\rm max}$ is in MeV, and $\alpha$ and $\beta$ are in $10^{-9}\,{\rm MeV}^{-3}$.
$q_{\rm max}$ $\alpha$ $\beta$ $M_{R},\Gamma_{R}]$ $|g_{\Omega(2012)\bar{K}\Xi(1530)}|$ $|g_{\Omega(2012)\eta\Omega}|$ $|g_{\Omega(2012)\bar{K}\Xi}|$ ${\cal R}^{\bar{K}\pi\Xi}_{\bar{K}\Xi}$
800 $175\pm 10$ $316\pm 22$ $2014.1\pm 1.0,3.4\pm 0.3]$ $2.07\pm 0.02$ $1.87\pm 0.07$ $0.21\pm 0.02$ $0.76\pm 0.15$
850 $174\pm 11$ $371\pm 28$ $2014.1\pm 1.0,3.4\pm 0.2]$ $2.09\pm 0.02$ $1.56\pm 0.05$ $0.20\pm 0.02$ $0.81\pm 0.14$
900 $157\pm 11$ $372\pm 30$ $2014.1\pm 1.0,3.4\pm 0.2]$ $2.10\pm 0.02$ $1.35\pm 0.05$ $0.20\pm 0.02$ $0.81\pm 0.14$
1000 $120\pm 9$ $328\pm 28$ $2014.2\pm 1.0,3.4\pm 0.2]$ $2.10\pm 0.03$ $1.07\pm 0.04$ $0.21\pm 0.02$ $0.81\pm 0.14$
1100 $91\pm 6$ $273\pm 24$ $2014.2\pm 1.0,3.4\pm 0.2]$ $2.09\pm 0.03$ $0.88\pm 0.04$ $0.20\pm 0.02$ $0.81\pm 0.14$
1200 $69\pm 5$ $219\pm 22$ $2014.2\pm 1.0,3.5\pm 0.3]$ $2.08\pm 0.03$ $0.73\pm 0.05$ $0.21\pm 0.03$ $0.78\pm 0.16$
The coupling of the $\Omega(2012)$ to each relevant channel $g_{i}$ ($i=\Omega(2012)\bar{K}\Xi(1530)$, $\Omega(2012)\eta\Omega$, $\Omega(2012)\bar{K}\Xi$) can be determined via the residues of the scattering amplitude $T_{ij}$ at the pole position, which reads,
$T_{ij}=\frac{g_{i}g_{j}}{\sqrt{s}-z_{R}},$ (9)
where $z_{R}=M_{R}-i\Gamma_{R}/2$ denotes the $\Omega(2012)$ pole position. The theoretical results for these coupling constants are also shown in Table 1 for each scenario of model parameters and $q_{\rm max}$. Considering the dominant radiative decay mechanism shown in Fig. 3, we need the coupling constant of $g_{\Omega(2012)\bar{K}\Xi(1530)}$, and in the numerical study we take
$g_{\Omega(2012)K^{-}\Xi(1530)^{0}}$ $=$ $g_{\Omega(2012)\bar{K}^{0}\Xi(1530)^{-}}$ (10)
$=$ $-\frac{1}{\sqrt{2}}g_{\Omega(2012)\bar{K}\Xi(1530)}.$
In the present notation, we employ the isospin doublets ($\bar{K}^{0}$, $K^{-}$) and ($\Xi(1530)^{0}$, $\Xi(1530)^{-}$). The resulting $I=0$ state of the $\Xi(1530)\bar{K}$ system is written as [37]
[…]
It is important to note that, in Ref. [38], only the mass and width of the $\Omega(2012)$ were fitted, while the upper limit $R^{\bar{K}\pi\Xi}_{\bar{K}\Xi}<0.119$ was imposed merely as a constraint. In the present work, however, the updated experimental value $R^{\bar{K}\pi\Xi}_{\bar{K}\Xi}=0.99\pm 0.26\pm 0.06$ is included as an additional data point in the fit. This change in the fitting inputs explains the differences between the theoretical parameters obtained here and those reported in Ref. [31]. Moreover, the larger absolute values of $\alpha$ and $\beta$ lead to a larger coupling constant $g_{\Omega(2012)\bar{K}\Xi(1530)}$ and a smaller $g_{\Omega(2012)\bar{K}\Xi}$, which in turn allows the new, larger experimental ratio ${\cal R}^{\bar{K}\pi\Xi}_{\bar{K}\Xi}$ to be reproduced within uncertainties. These results further support the molecular interpretation of the $\Omega(2012)$ state.
@@ II.2 The effective Lagrangians for the $\Xi(1530)\Xi\gamma$ and $\Omega\bar{K}\Xi$ interaction
$\mathcal{L}_{\Xi(1530)\Xi\gamma}$ $=$ $i\frac{eg_{1}}{2m_{N}}F^{\mu\nu}\bar{\Xi}\gamma_{\mu}\gamma_{5}\Xi(1530)_{\nu}$ (12)
$\!\!\!\!\!\!\!\!\!\!\!\!\!-\frac{eg_{2}}{(2m_{N})^{2}}F^{\mu\nu}(\partial_{\mu}\bar{\Xi})\gamma_{5}\Xi(1530)_{\nu}+H.c.\,$,
@@
$\mathcal{L}_{\Omega\bar{K}\Xi}=\frac{f_{\Omega\Xi\bar{K}}}{m_{K}}\bar{\Omega}_{v}\Xi\partial^{v}\bar{K}\,$, (13)
@@ II.3 The radiative decay $\Omega(2012)\to\gamma\Omega$
$\mathcal{M}_{\Omega(2012)\to\gamma\Omega}$ $=\int\frac{d^{4}q}{(2\pi)^{4}}\frac{\widetilde{A}F^{2}(q_{ex})}{q^{2}-m_{\Xi(1530)}^{2}+i\varepsilon}\times$ (14)
@@
where $P$, $q$, and $q_{ex}$ are the four-momenta of the $\Omega(2012)$, $\Xi(1530)$, and the exchanged hyperon $\Xi$, respectively. $m_{\Xi(1530)}$ is the mass of $\Xi(1530)$, and
@@
$\left.\frac{2q_{\beta}q_{\nu}}{3m_{\Xi(1530)}^{2}}+\frac{\gamma_{\beta}q_{\nu}-\gamma_{\nu}q_{\beta}}{3m_{\Xi(1530)}}\right],$
$A_{3}^{\nu}$ $=$ $\frac{g_{\Omega(2012)\bar{K}\Xi(1530)}}{\sqrt{2}}u^{\nu}(P,s_{\Omega(2012)}).$ (19)
Here, $k_{1}$ ($\lambda$) and $k_{2}$ ($s_{\Omega}$) are the four-momenta (spin polarization) of the final photon and $\Omega$ baryon. $u^{\mu}(P,s_{\Omega(2012)})$ is a Rarita-Schwinger spinor for the $\Omega(2012)$ resonance with momentum $P$ and spin polarization $s_{\Omega(2012)}$. The expression for the photon polarization vector $\varepsilon^{\mu}(k_{1},\lambda)$ are given by
@@
$\begin{gathered}P=(P^{0},0,0,0)=(M_{\Omega(2012)},0,0,0),\\ q=(q_{0},\left\lvert\vec{q}\right\rvert\sin\theta\cos\phi,\left\lvert\vec{q}\right\rvert\sin\theta\sin\phi,\left\lvert\vec{q}\right\rvert\cos\theta),\\ k_{1}=(E_{\gamma},0,0,-E_{\gamma}),k_{2}=(E_{\Omega},0,0,E_{\gamma}),\end{gathered}$ (27)
@@
$E_{\gamma/\Omega}=\frac{M_{\Omega(2012)}^{2}\mp m_{\Omega}^{2}}{2M_{\Omega(2012)}},$ (28)
where $M_{\Omega(2012)}$ is the mass of $\Omega(2012)$, $E_{\gamma}(E_{\Omega})$ is the energy of the final photon ($\Omega$). We first integrate out the $q_{0}$ component utilizing the residue theorem [47] and introduce a cutoff $\Lambda$ for the integration over $|\vec{q}|$  [^1] The triangle loop integral depends on the cutoff parameter $\Lambda$ from the form factor of the exchanged $\Xi$ in Eq. (26). We have therefore set the value of $\Lambda$ in this integral to be the same as that in the form factor. This treatment not only simplifies the calculation but also reduces the number of free parameters in the theory.,
@@
$D$ $=$ $8\omega_{\Xi(1530)}\omega_{K}\omega_{\Xi}(P^{0}+\omega_{\Xi(1530)}+\omega_{K})$ (30)
$\times(P^{0}-\omega_{\Xi(1530)}+\omega_{K}+i\epsilon)$
$\times(P^{0}-\omega_{\Xi(1530)}-\omega_{K}+i\epsilon)(E_{\gamma}+\omega_{\Xi(1530)}+\omega_{\Xi})$
$\times(E_{\gamma}-\omega_{\Xi(1530)}+\omega_{\Xi}+i\epsilon)(E_{\gamma}-\omega_{\Xi(1530)}-\omega_{\Xi}+i\epsilon)$
@@
$N_{1}$ $=$ $4\omega_{K}\omega_{\Xi}(P^{0}+\omega_{\Xi(1530)}+\omega_{K})(E_{\gamma}+\omega_{\Xi(1530)}+\omega_{\Xi})$
@@
$\times(E_{\Omega}-\omega_{K}-\omega_{\Xi})\widetilde{t}|_{\omega_{\Xi(1530)}},$
$N_{2}$ $=$ $4\omega_{\Xi(1530)}\omega_{\Xi}(P^{0}-\omega_{\Xi(1530)}+\omega_{K})(E_{\gamma}+\omega_{\Xi(1530)}+\omega_{\Xi})$
$\times(E_{\gamma}-\omega_{\Xi(1530)}+\omega_{\Xi})(E_{\gamma}-\omega_{\Xi(1530)}-\omega_{\Xi})$
@@
$N_{3}$ $=$ $4\omega_{\Xi(1530)}\omega_{K}(P^{0}+\omega_{\Xi(1530)}+\omega_{K})(P^{0}-\omega_{\Xi(1530)}+\omega_{K})$
$\times(P^{0}-\omega_{\Xi(1530)}-\omega_{K})(E_{\gamma}-\omega_{\Xi(1530)}-\omega_{\Xi})$
@@
where $\widetilde{t}|_{x}$ is the value of $\widetilde{t}$ obtained at $q^{0}=x$ with $\widetilde{t}=\widetilde{A}F^{2}(q_{ex})$, and
$\omega_{\Xi(1530)}$ $=\sqrt{|\vec{q}|^{2}+m_{\Xi(1530)}^{2}},~~~~~\omega_{K}=\sqrt{|\vec{q}|^{2}+m_{K}^{2}},$ (32)
@@
$\Gamma_{\Omega(2012)\to\gamma\Omega}=\frac{E_{\gamma}}{32\pi M_{\Omega(2012)}^{2}}\overline{\sum_{s_{\Omega(2012)}}}\sum_{\lambda,s_{\Omega}}|\mathcal{M}_{\Omega(2012)\to\gamma\Omega}|^{2},$ (33)